# Bibliography.bib

### v2

@@ Abstract.
then $\operatorname{vol}(M)\leqslant\operatorname{vol}(\mathbb{S}^{n})$, and $\pi_{1}(M)$ is finite. Moreover, the constant $\frac{n-1}{n-2}$ cannot be improved, and if $\operatorname{vol}(M)=\operatorname{vol}(\mathbb{S}^{n})$ holds, then $M\cong\mathbb{S}^{n}$. A sharp generalization of the Bonnet–Myers theorem is also shown under the same spectral condition.
@@ 1.1. Statement of the results
Observe that in Theorem 1, and Corollary 1 the function $u$ is assumed to be smooth. On the other hand, a natural condition that one usually encounters is the positivity of the Dirichlet energy
@@
    Let us emphasize the difference between Theorem 1 and Corollary 1: when $n>3$ and for the values $\gamma\in\big{[}\frac{4}{n-1},\frac{n-1}{n-2}\big{]}$, the manifold $M$ satisfying the main spectral condition (1) may not be compact (counterexamples include, for example, hyperbolic spaces). However, assuming that $M$ is compact, we have the volume and diameter bound, and the finiteness of $\pi_{1}(M)$, as a consequence of Theorem 1.
@@
Let us make the following connections and comparisons with the existing results.
@@
    In dimension $n=3$, Item 2 in Theorem 1 appeared in [CLMS, Theorem 5.1], and it has been relevant for the proof of the stable Bernstein Theorem in $\mathbb{R}^{5}$. The generalization to arbitrary dimensions requires a nontrivial change in the proof: see Section 1.2 for additional details.
@@
    Based on a Bochner formula, the following vanishing theorem is proved in [BourCarron]: if (1) holds with $\gamma=\frac{n-1}{n-2}$ and $\lambda=0$, then we have either $b_{1}(M)=0$, or $M$ splits as a warped product $(M,g)=(\mathbb{S}^{1}\times N,\mathrm{d}r^{2}+\eta(r)^{2}h)$, where $h$ is a metric on $N$ with $\operatorname{Ric}_{h}\geqslant 0$. By inspecting the computations in [BourCarron, p.107] we can also obtain that: for the positive case $\lambda>0$ the only possibility is $b_{1}(M)=0$, and for sub-critical values $\gamma<\frac{n-1}{n-2}$ (with $\lambda=0$), the result strengthens to that either $b_{1}(M)=0$, or $M$ splits as a direct product $(\mathbb{S}^{1}\times N,\mathrm{d}r^{2}+h)$.
@@ 1.2. Strategy of the proof
The same sharp inequality $I^{\prime\prime}I\leqslant-\frac{1}{2}(I^{\prime})^{2}-2\lambda$ is shown in [CLMS], which eventually leads to a sharp volume bound. In dimensions other than 3, however, the inequality obtained through this argument requires the narrower condition $\gamma<\frac{4}{n-1}$, and is not sharp.
@@
$I^{\prime\prime}I\leqslant-\frac{(I^{\prime})^{2}}{n-1}-(n-1)\lambda,$ (10)
and leads to the proof of the sharp volume bound (3).
@@
Added notes. Besides the application discussed in Theorem 2, we expect the results of this paper to be useful to deal with a variety of geometric problems. After this paper was made public, the main result Theorem 1, together with the strategy devised in [CLMS], has been exploited by L. Mazet to settle the stable Bernstein problem in $\mathbb{R}^{6}$ [Mazet]. Theorem 1 could pave the way to study the stable Bernstein problem in the remaining dimension $\mathbb{R}^{7}$. In addition, it would be interesting to understand the implications of our results to the study of the structure of manifolds (or limits of manifolds) with uniform Kato bounds on the Ricci curvature: see, e.g., [CarronMondelloTewodrose].
@@ 1.5. Acknowledgments
G.A. acknowledges the financial support of the Courant Institute, and the AMS-Simons Travel grant. K.X. would like to thank the hospitality of the Courant Institute, where part of this work was done during his visit. The authors are grateful to Chao Li, Xian-Tao Huang, Marco Pozzetta, and Alberto Roncoroni for interesting discussions and suggestions on the topic of the paper. The authors are grateful to Otis Chodosh for comments that led to the improvement of a previous version of Theorem 2, and to Jintian Zhu for comments that led to Remark 2.
@@ 3. Unequally weighted isoperimetric profile and volume bounds
For $\varphi\in C^{\infty}(M)$, let us consider a smooth family of sets $\{E_{t}\}_{t\in(-\varepsilon,\varepsilon)}$, such that $E_{0}=E$ and the variational vector field along $\partial E_{t}$ is $\varphi\nu$ (where $\nu$ denotes the outer unit normal of $\partial E_{t}$). Define $V(t):=V(E_{t})$ and $A(t):=A(E_{t})$; thus $V(0)=v_{0}$, $A(0)=I(v_{0})$. Let $H$ be the mean curvature of $\partial E$, and denote $\varphi_{\nu}:=\langle\nabla\varphi,\nu\rangle$, $u_{\nu}:=\langle\nabla u,\nu\rangle$ for brevity. We compute the first and second variations of the (weighted) volume at the initial time:
@@
Notice that when $\zeta=n\operatorname{vol}(\mathbb{B}^{n})=\operatorname{vol}(\mathbb{S}^{n})$, we exactly have $V_{\zeta}=\lambda^{-n/2}\operatorname{vol}(\mathbb{S}^{n})$.
@@ 4. Isoperimetry under $\mathrm{Ric}\geqslant 0$ and spectral $\mathrm{biRic}\geqslant n-2$
Let us present a proof of (80) under the non-collapsing assumption (79). With (79) available, we are free to use the results of [XingyuZhu]. Take a radial geodesic $\gamma$ issuing from $o$, and let $b$ be the associated Busemann function. Since $M$ has linear volume growth by Item 1 in Theorem 2, and since $\mathrm{Ric}\geqslant 0$, we can use [Sormani, Corollary 23] and infer $0\geqslant m:=\inf_{p\in M}b(p)>-\infty$. In addition, by [XingyuZhu, Proposition 2.6], we get that there is a constant $C>0$ such that $d(p,\gamma)\leqslant C$ for every $p\in M$. Hence we can freely apply [XingyuZhu, Proposition 2.10], and [XingyuZhu, Theorem 3.2] in what follows.
In particular, by [XingyuZhu, (2.16)] the function $t\mapsto|\partial^{*}\{b=t\}|$ is nondecreasing, and combining [XingyuZhu, (2.17)], and [XingyuZhu, Theorem 3.2(3)], we have
$\lim_{t\to\infty}|\partial^{*}\{b=t\}|=\lim_{v\to\infty}I_{M}(v).$ (81)
Moreover, from the definition of Busemann function, we have $b(x)\leqslant d(x,o)$ for every $x\in M$. Thus, by the coarea formula, for every $r>0$ we have
$\operatorname{vol}(B_{r}(o))\leqslant\operatorname{vol}(\{b<r\})=\int_{m}^{r}|\partial^{*}\{b=t\}|\,\mathrm{d}t\leqslant(r-m)|\partial^{*}\{b=r\}|.$
The latter, joined with the result proved in Item 2 in Theorem 2, and (81), implies (80) when the additional assumption $\inf_{p\in M}\operatorname{vol}(B_{1}(p))>0$ is in force.
@@ Appendix A The case $n\geqslant 8$
When trying to prove Theorem 1 in dimensions $n\geqslant 8$, one encounters the issue that the minimizer in Lemma 1 or Lemma 2 may contain singularities. In this section, we present a way to modify the second variation argument so that it is applicable to possibly singular minimizers. Similar arguments have also appeared in [Bayle04] and [BraySingular].
@@
For $\varphi\in C^{\infty}(M)$, consider a family of sets $\{E_{t}\}_{t\in(-\varepsilon,\varepsilon)}$, such that $E_{0}=E$ and the variational vector field along $\partial E_{t}$ is $\varphi\eta\nu$. This family is well-defined since $\eta$ is supported inside $\partial^{\text{reg}}E$. The variations of the area and the volume remain unchanged as in Lemma 2 (with each $\varphi$ replaced with $\varphi\eta$):

### v3

@@ Abstract.
then $\operatorname{vol}(M)\leqslant\operatorname{vol}(\mathbb{S}^{n})$, and $\pi_{1}(M)$ is finite. The constant $\frac{n-1}{n-2}$ cannot be improved, and if $\operatorname{vol}(M)=\operatorname{vol}(\mathbb{S}^{n})$ holds, then $M\cong\mathbb{S}^{n}$. A sharp generalization of the Bonnet–Myers theorem is also shown under the same spectral condition.
@@
Furthermore, the main result of this paper is applied in Mazet’s recent solution of the stable Bernstein problem in $\mathbb{R}^{6}$.
@@ 1.1. Statement of the results
Observe that in (1) and (4), the function $u$ is assumed to be smooth. On the other hand, a natural condition that one usually encounters is the positivity of the Dirichlet energy
@@
    Let us emphasize the difference between Theorem 1 and Corollary 1: when $n>3$ and $\gamma\in\big{[}\frac{4}{n-1},\frac{n-1}{n-2}\big{]}$, the manifold $M$ satisfying the main spectral condition (1) may not be compact (counterexamples include, for example, hyperbolic spaces). However, assuming that $M$ is compact, we have the volume and diameter bound, and the finiteness of $\pi_{1}(M)$, as a consequence of Theorem 1.
@@
When $n=3$, Item 2 in Theorem 1 has appeared in [CLMS, Theorem 5.1], as one step in the proof of the stable Bernstein Theorem in $\mathbb{R}^{5}$. After this paper was made public, the main Theorem 1 (with $n=4$) together with the strategy in [CLMS] has been exploited by L. Mazet to resolve the stable Bernstein problem in $\mathbb{R}^{6}$. We remark that the range of $\gamma$ appearing in Theorem 1 is crucially related to the dimension restriction for the strategies used in [CLMS, Mazet].
Theorem 2 ([Mazet, Theorem 1.1]).
Let $M^{5}\hookrightarrow\mathbb{R}^{6}$ be an immersed, complete, connected, two-sided, stable minimal hypersurface. Then $M$ is a Euclidean hyperplane.
Let us make some further connections and comparisons with the existing results.
@@
    Based on a Bochner formula from [LiWang], the following vanishing theorem is proved in [BourCarron]: if (1) holds with $\gamma=\frac{n-1}{n-2}$ and $\lambda=0$, then we have either $b_{1}(M)=0$, or $M$ is isometric to a quotient of the warped product (×$(\mathbb{R}\times N,\mathrm{d}r^{2}+\eta(r)^{2}h)$, where $h$ is a metric on $N$ with $\operatorname{Ric}_{h}\geqslant 0$. By inspecting the computations in [BourCarron, p.107] we can also obtain that: for the positive case $\lambda>0$ the only possibility is $b_{1}(M)=0$, and for sub-critical values $\gamma<\frac{n-1}{n-2}$ (with $\lambda=0$), the result strengthens to that either $b_{1}(M)=0$, or $M$ is a quotient of the direct product (×$(\mathbb{R}\times N,\mathrm{d}r^{2}+h)$.
@@
  8. (viii)
    After we posted our paper, [JiaLi, ChuHao] has proved similar results for the case of manifolds with boundary and for Bakry-Émery curvature conditions, respectively.
@@ 1.2. Strategy of the proof
The same sharp inequality $I^{\prime\prime}I\leqslant-\frac{1}{2}(I^{\prime})^{2}-2\lambda$ is shown in [CLMS], which eventually leads to a sharp volume bound. In other dimensions, however, the inequality obtained through this argument requires the condition $\gamma<\frac{4}{n-1}$ (which is strictly narrower than $\gamma\leqslant\frac{n-1}{n-2}$ when $n>3$), and the volume bound is not sharp.
@@
$I^{\prime\prime}I\leqslant-\frac{(I^{\prime})^{2}}{n-1}-(n-1)\lambda.$ (10)
This leads to the proof of the sharp volume bound (3).
@@
In the very recent paper [ZhuZhou], the authors answer negatively to 1 when $n\in\{6,7\}$ by refining the $\mu$-bubble argument above: see [ZhuZhou, Theorem 1.5]. 1 remains open when $n\geqslant 8$.
@@ 1.5. Acknowledgments
G.A. acknowledges the financial support of the Courant Institute, and the AMS-Simons Travel grant. K.X. would like to thank the hospitality of the Courant Institute, where part of this work was done during his visit. The authors are grateful to Chao Li, Xian-Tao Huang, Laurent Mazet, Luciano Mari, Marco Pozzetta, and Alberto Roncoroni for interesting discussions and suggestions on the topic of the paper. The authors are grateful to Otis Chodosh for comments that led to the improvement of a previous version of Theorem 3, and to Jintian Zhu for comments that led to Remark 2.
@@ 3. Unequally weighted isoperimetric profile and volume bounds
For $\varphi\in C^{\infty}(M)$, let us consider a smooth family of sets $\{E_{t}\}_{t\in(-\varepsilon,\varepsilon)}$, such that $E_{0}=E$, the variational vector field $X_{t}$ along $\partial E_{t}$ at $t=0$ is $\varphi\nu$ (where $\nu$ denotes the outer unit normal of $\partial E_{t}$), and $\nabla_{X_{t}}X_{t}=(\varphi\varphi_{\nu})\nu$ at $t=0$. Define $V(t):=V(E_{t})$ and $A(t):=A(E_{t})$; thus $V(0)=v_{0}$, $A(0)=I(v_{0})$. Let $H$ be the mean curvature of $\partial E$, and denote $\varphi_{\nu}:=\langle\nabla\varphi,\nu\rangle$, $u_{\nu}:=\langle\nabla u,\nu\rangle$ for brevity. We compute the first and second variations of the (weighted) volume at the initial time:
@@
Notice that when $\zeta=n\operatorname{vol}(\mathbb{B}^{n})=\operatorname{vol}(\mathbb{S}^{n-1})$, we exactly have $V_{\zeta}=\lambda^{-n/2}\operatorname{vol}(\mathbb{S}^{n})$.
@@ 4. Isoperimetry under $\mathrm{Ric}\geqslant 0$ and spectral $\mathrm{biRic}\geqslant n-2$
Let us present how to deduce (80) from the results in [XingyuZhu] under the non-collapsing assumption (79). With (79) available, we are free to use the results of [XingyuZhu]. Notice that $M$ has linear volume growth by Item 1 in Theorem 3. In particular, by combining [XingyuZhu, Corollary 4.3], and Item 2 in Theorem 3 we conclude
$\limsup_{r\to+\infty}\frac{\operatorname{vol}(B_{r}(o))}{r}=\lim_{v\to\infty}I_{M}(v)\leqslant\operatorname{vol}(\mathbb{S}^{n-1}).$
@@ Appendix A The case $n\geqslant 8$
When trying to prove Theorem 1 in dimensions $n\geqslant 8$, one encounters the issue that the minimizer in Lemma 1 or Lemma 2 may contain singularities. In this section, we present a way to modify the second variation argument so that it is applicable to possibly singular minimizers. Similar arguments have also appeared in [Bayle04, BraySingular, JonathanZhu].
@@
For $\varphi\in C^{\infty}(M)$, let us consider a smooth family of sets $\{E_{t}\}_{t\in(-\varepsilon,\varepsilon)}$, such that $E_{0}=E$, the variational vector field $X_{t}$ along $\partial E_{t}$ at $t=0$ is $\varphi\eta\nu$ (where $\nu$ denotes the outer unit normal of $\partial E_{t}$), and $\nabla_{X_{t}}X_{t}=(\varphi\eta(\varphi\eta)_{\nu})\nu$ at $t=0$. This family is well-defined since $\eta$ is supported inside $\partial^{\text{reg}}E$. The variations of the area and the volume remain unchanged as in Lemma 2 (with each $\varphi$ replaced with $\varphi\eta$):